# Markovian dynamics for a quantum/classical system and quantum trajectories

### v1

@@ 2.2 Stochastic Schrödinger equations
We introduce now the SDE which determines the dynamics of the quantum component and the interaction of the two components. The coefficients involved in such an equation are given in the following assumption; we consider only bounded operators on the Hilbert space of the quantum component to avoid analytical complications.
@@
$H(x)^{\dagger}=H(x),\qquad\sup_{x\in\mathbb{R}^{s}}\left\|H(x)\right\|<+\infty,\quad\sup_{x\in\mathbb{R}^{s}}\left\|L_{k}(x)\right\|<+\infty,\quad\sup_{x\in\mathbb{R}^{s},\;u\in\mathscr{U}}\left\|J(x,u)\right\|<+\infty,$ (6a)
$R(x):=\int_{\mathscr{U}}J(x,u)^{\dagger}J(x,u)\,\nu(\mathrm{d}u)\in\mathscr{B}(\mathscr{H}),\qquad\sup_{x\in\mathbb{R}^{s}}\left\|R(x)\right\|<+\infty;$ (6b)
the integral is defined in the weak topology of $\mathscr{B}(\mathscr{H})$. margin: $R(x)$
Then, we define
margin: $K(x)$
$K(x)=\mathrm{i}H(x)+\frac{1}{2}\sum_{k=1}^{d}L_{k}(x)^{\dagger}L_{k}(x)+\frac{1}{2}\,R(x),\qquad K_{t}(\omega)=K\big{(}X(t_{-};\omega)\big{)},\qquad H_{t}(\omega)=H\big{(}X(t_{-};\omega)\big{)},$ (7a)
$L_{kt}(\omega)=L_{k}\big{(}X(t_{-};\omega)\big{)},\qquad J_{t}(u;\omega)=J\big{(}X(t_{-};\omega),u\big{)},\qquad R_{t}(\omega)=R\big{(}X(t_{-};\omega)\big{)}.$ (7b)
Let us stress that these coefficients depend on the solution of (2) and, so, on the initial condition $X(0)$. Finally, we introduce the linear stochastic Schrödinger equation (SSE) in $\mathscr{H}$ [31, (2.3)] margin: SSE
$\mathrm{d}\psi_{t}=-K_{t}\psi_{t_{-}}\mathrm{d}t+\sum_{k=1}^{d}L_{kt}\psi_{t_{-}}\,\mathrm{d}W_{k}(t)+\int_{\mathscr{U}}J_{t}(u)\psi_{t_{-}}\,\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t),\qquad\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{0}\right\|^{2}\big{]}<+\infty;$ (8)
@@
Let us note that the assumption of “no memory” in the dynamics of the composed system is implicitly contained in the fact that the coefficients in (2) and (8) depend on the past only through $X(t_{-})$ and that the involved noises are Wiener and Poisson processes, which have increments independent of the past. Examples without these hypotheses are considered in [31]. However, with respect to what is done in that reference, the space $\mathscr{U}$ involved in the point process is slightly more general; similarly, the structure of the classical process $X(t)$ has a jump contribution more general than in [31, (4.32)]. In any case, the proofs of that article apply to this case too and we can follow that construction.
@@
Up to $\mathbb{Q}$-equivalence, Eq. (8) admits a unique solution $\psi_{t}$, $t\in\mathbb{R}^{+}$, which is an $\mathscr{H}$-valued semimartingale [37, Sec. II.1]. The process $\left\|\psi_{t}\right\|^{2}$ is a non-negative martingale [37, Sec. I.2], such that margin: $\psi_{t},\ \widehat{\psi}_{t}$
$\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{t}\right\|^{2}\big{]}=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{0}\right\|^{2}\big{]}<+\infty,\qquad\mathrm{d}\left\|\psi_{t}\right\|^{2}=\left\|\psi_{t_{-}}\right\|^{2}\,\mathrm{d}Z_{t},$ (9a)
$Z_{t}=\sum_{k=1}^{d}\int_{(0,t]}m_{ks}\,\mathrm{d}W_{k}(s)+\int_{(0,t]}\int_{\mathscr{U}}\left(I_{s}(u)-1\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}s),\qquad m_{kt}=2\mathrm{\,Re\,}\langle\widehat{\psi}_{t_{-}}\big{|}L_{kt}\widehat{\psi}_{t_{-}}\rangle,$ (9b)
[…]
@@
The non-linear SSE.
The process $\widehat{\psi}_{t}$, $0\leq t\leq T$, satisfies a non-linear stochastic differential equation, the non-linear SSE, under the new probability $\left\|\psi_{T}(\omega)\right\|^{2}\mathbb{Q}(\mathrm{d}\omega)$, see [31, Sec. 2.2]. A first example of SSE goes back to [16].
@@ 2.3 Stochastic master equations
We introduce now the SDE in $\mathscr{T}(\mathscr{H})$ associated to the linear SSE (8).
@@
By the formal rules of stochastic calculus, we get the linear stochastic master equation (SME) for $\sigma_{t}$: margin: SME a()for-$\forall a\in\mathscr{B}(\mathscr{H})$, margin: $\mathcal{L}_{t}$ margin: $\mathcal{L}(x)$
@@
$\mathcal{L}_{t}=\mathcal{L}\big{(}X(t_{-})\big{)},$ (13a)
$\mathcal{L}(x)[a]=\mathrm{i}[H(x),a]+\sum_{k=1}^{d}\left(L_{k}(x)^{\dagger}aL_{k}(x)-\frac{1}{2}\left\{L_{k}(x)^{\dagger}L_{k}(x),a\right\}\right)\\ {}+\int_{\mathscr{U}}\left(J(x,u)^{\dagger}aJ(x,u)-\frac{1}{2}\left\{J(x,u)^{\dagger}J(x,u),a\right\}\right)\nu(\mathrm{d}u).$ (13b)
@@
$\sigma_{t}\geq 0,\qquad p_{t}:=\langle\sigma_{t},\mathds{1}\rangle\equiv\operatorname{Tr}\{\sigma_{t}\}\geq 0,\qquad\operatorname{\mathbb{E}}_{\mathbb{Q}}[p_{t}]=1,$ (14a)
$\mathrm{d}p_{t}=p_{t_{-}}\biggl{(}\sum_{k=1}^{d}m_{k}(t)\,\mathrm{d}W_{k}(t)+\int_{\mathscr{U}}\left(I_{t}(u)-1\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)\biggr{)},$ (14b)
$m_{k}(t)=2\mathrm{\,Re\,}\langle\widehat{\sigma}_{t_{-}},L_{kt}\rangle,\qquad I_{t}(u)=\big{\langle}\widehat{\sigma}_{t_{-}},\big{(}J_{t}(u)^{\dagger}+\mathds{1}\big{)}\big{(}J_{t}(u)+\mathds{1}\big{)}\big{\rangle},$ (14c)
[…]
@@
Again Assumptions 1, 3, and the properties of $X(t)$ imply the validity of Properties [31, 3.1, 3.3]. Then, Propositions 3.2 and 3.4 of [31] give the thesis of the theorem. ∎
By definition $\widehat{\sigma}_{t}(\omega)\in\mathscr{S}(\mathscr{H})$, i.e. $\widehat{\sigma}_{t}$ is a random quantum state. The random quantities $m_{k}(t)$ and $I_{t}(u)$ are defined by (7b), (14c), so as to be $\mathscr{F}_{t_{-}}$-measurable (predictable processes [33, Sec. 3]). By construction, if the initial state $\widehat{\sigma}_{0}$ is a pure state, then $\widehat{\sigma}_{t}(\omega)$ too is a pure state.
2.3.1 The physical probability and the classical component
As in [31, Sec. 3.1], we use $p_{t}$ (14a) as a probability density and we define the new probability $\mathbb{P}$ on the $\sigma$-algebra $\mathscr{F}=\bigvee_{t\geq 0}\mathscr{F}_{t}$ by
@@ 2.3.1 The physical probability and the classical component
for a similar construction of a probability measure see [33, Sec. 30.2].
@@
The probability $\mathbb{P}$, margin: $\mathbb{P}$ defined by (15), represents the physical probability of the hybrid system. By (2.2) and (12), the physical probability depends on the dynamics of both the classical component and the quantum one. Moreover, the probability $\mathbb{P}$ depends on the initial conditions $X(0)$ and $\sigma_{0}$ used in solving equations (2) and (12).
@@
By the so called Girsanov transformation and generalizations [37, Sec. III.8], we have that, under the probability $\mathbb{P}$, the processes
@@
are independent standard Wiener processes adapted to the filtration $\{\mathscr{F}_{t},\;t\geq 0\}$; moreover, $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson process, but it becomes a point process with compensator $I_{t}(u)\nu(\mathrm{d}u)\mathrm{d}t$ [31, Prop. 2.5, Remarks 2.6, 3.5]. We also introduce the notation
@@
According to [33, Sec. 31], $\widehat{\Pi}(\mathrm{d}u,\mathrm{d}t)$ is a white random measure.
@@
$\mathrm{d}X_{i}(t)=\left(c_{i}\big{(}X(t_{-})\big{)}+C^{(1)}_{i}(t)+C^{(2)}_{i}(t)\right)\mathrm{d}t+\sum_{k=1}^{d}a_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}\widehat{W}_{k}(t)\\ {}+\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\widehat{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\Pi(\mathrm{d}u,\mathrm{d}t),$ (18a)
$C^{(1)}_{i}(t)=\sum_{k=1}^{d}a_{i}^{k}\big{(}X(t_{-})\big{)}m_{k}(t),\qquad C^{(2)}_{i}(t)=\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\left(I_{t}(u)-1\right)\nu(\mathrm{d}u).$ (18b)
Here, $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson measure and the processes $m_{k}(t)$ and $I_{t}(u)$ depend on the whole past through the quantum state; in particular, the dynamics (2.3.1) of the classical component depends on the behaviour of the quantum one.
From Eqs. (2.3.1) we have that the classical component can extract some information from the quantum one only when not all the terms $C^{(1)}(t)$, $C^{(2)}(t)$, $I_{t}(u)$ vanish and they effectively depend on the quantum state. So, firstly we need $a_{i}^{k}(x)$ and $g_{i}(x,u)$ not all identically vanishing and this means that the diffusive and the jump terms are not both vanishing: some dissipation must be present in the dynamics (18a) of the classical component. Moreover, to have dependence on the quantum state at least some of the operators $L_{k}(x)$ and $J(x,u)$ must be not reducible to something proportional to the identity operator; this means that the Liouville operator (13b) must contain a dissipative contribution. As a consequence, also in this class of models of hybrid dynamics, we have that, in order to extract information from the quantum component, we need a dissipative dynamics for the quantum system. Moreover, the classical system acting as a measuring apparatus, as it receives this information, has necessarily a dissipative dynamics, because of the probabilistic nature of any quantum information.
@@
2.3.2 The non-linear SME and the conditional states
By the rules of stochastic calculus we can obtain the non-linear SME satisfied by the normalized random states $\widehat{\sigma}_{t}$ defined in (14d):
@@ 2.3.2 The non-linear SME and the conditional states
$\mathsf{E}_{t}$ is the random subset of $\mathscr{U}$, defined by $\mathsf{E}_{t}(\omega)=\{u\in\mathscr{U}:I_{t}(u;\omega)\neq 0\}$. Equation (19) follows from [31, Remark 3.6].
@@
$\rho_{t}:=\operatorname{\mathbb{E}}_{\mathbb{P}}[\widehat{\sigma}_{t}|\mathscr{E}_{t}].$ (20)
When the two filtrations coincide, $\mathscr{E}_{t}=\mathscr{F}_{t}$, the conditional states satisfy the non-linear SME (19); in general, the possibility of having a closed evolution equation for the conditional states depends on the structure of the coefficients; examples are given in [31, Sec. 4].
@@
where the random Liouvillian ${\mathcal{L}_{t}}_{*}$ is the pre-adjoint of the operator defined by (2.3). When the coefficients in the definition of $\mathcal{L}(x)$ do not depend on $x$, we get that ${\mathcal{L}_{t}}_{*}$ is non random and the a priori state satisfies a closed quantum master equation. In general, equation (22) is not closed and the reduced dynamics of the quantum component is not Markovian. The $x$-dependence in the Louville operator (13b) gives the dependence of the quantum dynamics on the classical component and it controls the transfer of information from the classical system to the quantum one. This $x$-dependence represents a feedback of the classical system on the quantum one; moreover, it could appear in the Hamiltonian term alone: the information transfer from the classical to the quantum component is not related to dissipativity.

### v2

@@ 2.2 Stochastic Schrödinger equations
We introduce now the stochastic equation which determines the dynamics of the quantum component and the interaction of the two components. The coefficients involved in such an equation are given in the following assumption; to avoid analytical complications, we consider only bounded operators on the Hilbert space of the quantum component.
@@
$\begin{split}H(x)^{\dagger}=H(x),\qquad&\sup_{x\in\mathbb{R}^{s}}\left\|H(x)\right\|<+\infty,\qquad\sup_{x\in\mathbb{R}^{s}}\left\|L_{k}(x)\right\|<+\infty,\qquad\sup_{x\in\mathbb{R}^{s},\;u\in\mathscr{U}}\left\|J(x,u)\right\|<+\infty,\\ &D(x):=\int_{\mathscr{U}}J(x,u)^{\dagger}J(x,u)\,\nu(\mathrm{d}u)\in\mathscr{B}(\mathscr{H}),\qquad\sup_{x\in\mathbb{R}^{s}}\left\|D(x)\right\|<+\infty;\end{split}$ (4)
the integral is defined in the weak topology of $\mathscr{B}(\mathscr{H})$. margin: $D(x)$
These operators become operator-valued random variables when composed with $X(t_{-})$. Indeed, we define, for $x\in\mathbb{R}^{s}$ and $\omega\in\Omega$, margin: $K(x)$
$K(x)=\mathrm{i}H(x)+\frac{1}{2}\sum_{k=1}^{d}L_{k}(x)^{\dagger}L_{k}(x)+\frac{1}{2}\,D(x),$ (5)
$\begin{split}K_{t}(\omega)=K\big{(}X(t_{-};\omega)\big{)},\qquad&H_{t}(\omega)=H\big{(}X(t_{-};\omega)\big{)},\\ L_{kt}(\omega)=L_{k}\big{(}X(t_{-};\omega)\big{)},\qquad J_{t}(u;\omega)=&J\big{(}X(t_{-};\omega),u\big{)},\qquad D_{t}(\omega)=D\big{(}X(t_{-};\omega)\big{)}.\end{split}$ (6)
We have inserted $X(t_{-})$ in order to have predictable operator-valued processes, as was for the coefficients in Eq. (2). Let us stress that these coefficients depend on the solution of (2) and, so, on the initial condition $X(0)$.
Finally, we introduce the linear stochastic Schrödinger equation (SSE) in $\mathscr{H}$ [36, (2.3)] margin: SSE
$\mathrm{d}\psi_{t}=-K_{t}\psi_{t_{-}}\mathrm{d}t+\sum_{k=1}^{d}L_{kt}\psi_{t_{-}}\,\mathrm{d}W_{k}(t)+\int_{\mathscr{U}}J_{t}(u)\psi_{t_{-}}\,\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t),\quad\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{0}\right\|^{2}\big{]}<+\infty;$ (7)
@@
Let us note that the assumption of “no memory” in the dynamics of the composed system is implicitly contained in the fact that the coefficients in (2) and (7) depend on the past only through $X(t_{-})$ and that the involved noises are Wiener and Poisson processes, which have increments independent of the past. Examples without these hypotheses are considered in [36]. However, with respect to what is done in that reference, the space $\mathscr{U}$ involved in the point process is slightly more general; similarly, the structure of the classical process $X(t)$ has a jump contribution more general than in [36, (4.32)]. In any case, the proofs of that article apply to this case too and we can follow that construction.
@@
Up to $\mathbb{Q}$-equivalence, Eq. (7) admits a unique solution $\psi_{t}$, $t\in\mathbb{R}^{+}$, which is an $\mathscr{H}$-valued semimartingale [42, Sec. II.1]. The process $\left\|\psi_{t}\right\|^{2}$ is a non-negative martingale [42, Sec. I.2], such that margin: $\psi_{t},\ \widehat{\psi}_{t}$
$\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{t}\right\|^{2}\big{]}=\operatorname{\mathbb{E}}_{\mathbb{Q}}\big{[}\left\|\psi_{0}\right\|^{2}\big{]}<+\infty,\qquad\mathrm{d}\left\|\psi_{t}\right\|^{2}=\left\|\psi_{t_{-}}\right\|^{2}\,\mathrm{d}Z_{t},$
$Z_{t}=\sum_{k=1}^{d}\int_{(0,t]}m_{ks}\,\mathrm{d}W_{k}(s)+\int_{(0,t]}\int_{\mathscr{U}}\left(I_{s}(u)-1\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}s),\quad m_{kt}=2\mathrm{\,Re\,}\langle\widehat{\psi}_{t_{-}}\big{|}L_{kt}\widehat{\psi}_{t_{-}}\rangle,$
$I_{t}(u)=\left\|\big{(}J_{t}(u)+\mathds{1}\big{)}\widehat{\psi}_{t_{-}}\right\|^{2},\qquad\widehat{\psi}_{t}(\omega)=\begin{cases}\psi_{t}(\omega)/\left\|\psi_{t}(\omega)\right\|&\text{if }\ \left\|\psi_{t}(\omega)\right\|\neq 0,\\ v\ {\rm(fixed\ unit\ vector)}&\text{if }\ \left\|\psi_{t}(\omega)\right\|=0.\end{cases}$
@@
Remark 1 (The non-linear SSE).
The process $\widehat{\psi}_{t}$, $0\leq t\leq T$, satisfies a non-linear stochastic differential equation, the non-linear SSE, under the new probability $\left\|\psi_{T}(\omega)\right\|^{2}\mathbb{Q}(\mathrm{d}\omega)$, see [36, Sec. 2.2]. A first example of SSE goes back to [16].
@@ 2.3 Stochastic master equations
We introduce now the stochastic differential equation in $\mathscr{T}(\mathscr{H})$ associated to the linear SSE (7).
@@
By the formal rules of stochastic calculus, we get the linear stochastic master equation margin: SME (SME) for $\sigma_{t}$: a()for-$\forall a\in\mathscr{B}(\mathscr{H})$, margin: $\mathcal{L}_{t}$ margin: $\mathcal{L}(x)$
@@
$\mathcal{L}_{t}=\mathcal{L}\big{(}X(t_{-})\big{)},$ (11)
$\mathcal{L}(x)[a]=\mathrm{i}[H(x),a]+\sum_{k=1}^{d}\left(L_{k}(x)^{\dagger}aL_{k}(x)-\frac{1}{2}\left\{L_{k}(x)^{\dagger}L_{k}(x),a\right\}\right)\\ {}+\int_{\mathscr{U}}\left(J(x,u)^{\dagger}aJ(x,u)-\frac{1}{2}\left\{J(x,u)^{\dagger}J(x,u),a\right\}\right)\nu(\mathrm{d}u).$ (12)
Let us stress that the SME (10) is linear once the process $X(t)$, solving (2), as been fixed. Instead, if we consider (2) and (10) as a system of SDEs, we have the non-linearity introduced by (2).
@@
$\sigma_{t}\geq 0,\qquad p_{t}:=\langle\sigma_{t},\mathds{1}\rangle\equiv\operatorname{Tr}\{\sigma_{t}\}\geq 0,\qquad\operatorname{\mathbb{E}}_{\mathbb{Q}}[p_{t}]=1,$ (13)
$\mathrm{d}p_{t}=p_{t_{-}}\biggl{(}\sum_{k=1}^{d}m_{k}(t)\,\mathrm{d}W_{k}(t)+\int_{\mathscr{U}}\left(I_{t}(u)-1\right)\widetilde{\Pi}(\mathrm{d}u,\mathrm{d}t)\biggr{)},$ (14)
$m_{k}(t)=2\mathrm{\,Re\,}\langle\widehat{\sigma}_{t_{-}},L_{kt}\rangle,\qquad I_{t}(u)=\big{\langle}\widehat{\sigma}_{t_{-}},\big{(}J_{t}(u)^{\dagger}+\mathds{1}\big{)}\big{(}J_{t}(u)+\mathds{1}\big{)}\big{\rangle},$ (15)
$\widehat{\sigma}_{t}(\omega)=\begin{cases}\sigma_{t}(\omega)/p_{t}(\omega)&\text{if }\ p_{t}(\omega)\neq 0,\\ \sigma_{v}\ {\rm(fixed\ statistical\ operator)}&\text{if }\ p_{t}(\omega)=0.\end{cases}$ (16)
@@
Again Assumptions 1, 3, and the properties of $X(t)$ imply the validity of Properties 3.1 and 3.3 of [36]. Then, Propositions 3.2 and 3.4 of [36] give the thesis of the theorem. ∎
By definition $\widehat{\sigma}_{t}(\omega)\in\mathscr{S}(\mathscr{H})$, i.e. $\widehat{\sigma}_{t}$ is a random quantum state. The random quantities $m_{k}(t)$ and $I_{t}(u)$ are defined by (6), (15), so as to be $\mathscr{F}_{t_{-}}$-measurable (predictable processes [38, Sec. 3]). By construction, if the initial state $\widehat{\sigma}_{0}$ is a pure state, then $\widehat{\sigma}_{t}(\omega)$ too is a pure state.
2.3.1 The physical probability and the classical component.
As in [36, Sec. 3.1], we use $p_{t}$ (13) as a probability density and we define the new probability $\mathbb{P}$ on the $\sigma$-algebra $\mathscr{F}=\bigvee_{t\geq 0}\mathscr{F}_{t}$ by
@@
for a similar construction of a probability measure see [38, Sec. 30.2].
Note that $p_{t}$ (the trace of $\sigma_{t}$) is a probability density with respect to $\mathbb{Q}$ and that it gives the probability $\mathbb{P}$ on $\mathscr{F}_{t}$, which contains the events related to the time interval $(0,t]$. By (14), $p_{t}$ is a martingale and this implies the consistency of the probabilities constructed in different time intervals: for $0<s<t$, $\mathsf{E}\in\mathscr{F}_{s}$, one has $\int_{\mathsf{E}}p_{t}(\omega)\mathbb{Q}(\mathrm{d}\omega)=\int_{\mathsf{E}}p_{s}(\omega)\mathbb{Q}(\mathrm{d}\omega)$. This allows to extend $\mathbb{P}$ to the whole $\mathscr{F}$.
@@
The probability $\mathbb{P}$, margin: $\mathbb{P}$ defined by (17), represents the physical probability of the hybrid system. By (6) and (10), the physical probability depends on the dynamics of both the classical component and the quantum one. Moreover, the probability $\mathbb{P}$ depends on the initial conditions $X(0)$ and $\sigma_{0}$ used in solving equations (2) and (10).
@@
By the so called Girsanov transformation and generalizations [42, Sec. III.8], we have that, under the probability $\mathbb{P}$, the processes
@@
are independent standard Wiener processes adapted to the filtration $\{\mathscr{F}_{t},\;t\geq 0\}$; moreover, $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson process, but it becomes a point process with compensator $I_{t}(u)\nu(\mathrm{d}u)\mathrm{d}t$ [36, Prop. 2.5, Remarks 2.6, 3.5]. We also introduce the notation
@@
According to [38, Sec. 31], $\widehat{\Pi}(\mathrm{d}u,\mathrm{d}t)$ is a white random measure.
@@
$\mathrm{d}X_{i}(t)=\left(c_{i}\big{(}X(t_{-})\big{)}+C^{(1)}_{i}(t)+C^{(2)}_{i}(t)\right)\mathrm{d}t+\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}\mathrm{d}\widehat{W}_{k}(t)\\ {}+\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\widehat{\Pi}(\mathrm{d}u,\mathrm{d}t)+\int_{\mathscr{U}\setminus\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\Pi(\mathrm{d}u,\mathrm{d}t),$ (20)
$C^{(1)}_{i}(t)=\sum_{k=1}^{d}b_{i}^{k}\big{(}X(t_{-})\big{)}m_{k}(t),\qquad C^{(2)}_{i}(t)=\int_{\mathsf{U}_{0}}g_{i}\big{(}X(t_{-}),u\big{)}\left(I_{t}(u)-1\right)\nu(\mathrm{d}u).$ (21)
Here, $\Pi(\mathrm{d}u,\mathrm{d}t)$ is no more a Poisson measure and the processes $m_{k}(t)$ and $I_{t}(u)$ depend on the whole past through the quantum state; in particular, the dynamics (20) of the classical component depends on the behaviour of the quantum one.
From Eqs. (20), (21) we have that the classical component can extract some information from the quantum one only when not all the terms $C^{(1)}(t)$, $C^{(2)}(t)$, $I_{t}(u)$ vanish and they effectively depend on the quantum state. So, firstly we need $b_{i}^{k}(x)$ and $g_{i}(x,u)$ not all identically vanishing and this means that the diffusive and the jump terms are not both vanishing: some dissipation must be present in the dynamics (20) of the classical component. Moreover, to have dependence on the quantum state at least some of the operators $L_{k}(x)$ and $J(x,u)$ must be not reducible to something proportional to the identity operator; this means that the Liouville operator (12) must contain a dissipative contribution. As a consequence, also in this class of models of hybrid dynamics, we have that, in order to extract information from the quantum component, we need a dissipative dynamics for the quantum system. Furthermore, the classical system acting as a measuring apparatus, as it receives this information, has necessarily a dissipative dynamics, because of the probabilistic nature of any quantum information.
@@
2.3.2 The non-linear SME and the conditional states.
By the rules of stochastic calculus we can obtain the non-linear SME satisfied by the normalized random states $\widehat{\sigma}_{t}$ defined in (16):
@@
$\mathsf{E}_{t}$ is the random subset of $\mathscr{U}$, defined by $\mathsf{E}_{t}(\omega)=\{u\in\mathscr{U}:I_{t}(u;\omega)\neq 0\}$. Equation (22) follows from [36, Remark 3.6].
The two coupled SDEs (20) and (22) have been obtained by construction. This proves that these equations admit a solution and that it exists a probability $\mathbb{P}$ under which the processes $\widehat{W}_{k}$ are independent Wiener processes and $\Pi(\mathrm{d}u,\mathrm{d}t)$ is a point measure with random compensator $I_{t}(u)\nu(\mathrm{d}u)\mathrm{d}t$; the uniqueness remains open. In a simplified case, without jump contributions, uniqueness has been proved in [23, Theor. 5.2].
Remark 4 (Conditional state).
@@
$\rho_{t}:=\operatorname{\mathbb{E}}_{\mathbb{P}}[\widehat{\sigma}_{t}|\mathscr{E}_{t}]\equiv\operatorname{\mathbb{E}}_{\mathbb{P}}[\widehat{\sigma}_{t}|\mathscr{E}],\qquad\mathscr{E}=\bigvee_{t\geq 0}\mathscr{E}_{t}.$ (23)
When the two filtrations coincide, $\mathscr{E}_{t}=\mathscr{F}_{t}$, the conditional states satisfy the non-linear SME (22). In general, the possibility of having a closed evolution equation for the conditional states depends on the structure of the coefficients; examples are given in [36, Sec. 4]. When not all the components $X_{1}(t),\ldots,X_{s}(t)$ of the classical system are observed, one has to take as $\{\mathscr{E}_{t}\}$ the filtration generated by the observed components of $X(t)$; this is a case of partial observation and it opens to the application of all the notions and results of classical filtering theory [49], which indeed is at the origin of the idea of observations in continuous time in quantum theories [16].
Remark 5 (Mean state).
@@
where the random Liouvillian ${\mathcal{L}_{t}}_{*}$ is the pre-adjoint of the operator defined by (11). When the coefficients in the definition of $\mathcal{L}(x)$ do not depend on $x$, we get that ${\mathcal{L}_{t}}_{*}$ is non random and the a priori state satisfies a closed quantum master equation. In general, equation (25) is not closed and the reduced dynamics of the quantum component is not Markovian. The $x$-dependence in the Liouville operator (12) gives the dependence of the quantum dynamics on the classical component and it controls the transfer of information from the classical system to the quantum one. This $x$-dependence represents a feedback of the classical system on the quantum one; moreover, it could appear in the Hamiltonian term alone: the information transfer from the classical to the quantum component is not related to dissipation.
The notions of conditional and mean states are typical of the approach to hybrid systems based on measurement and filtering [16, 22, 20, 23, 24], but the idea of considering the full dynamics and its reduced (or marginal) version appears also in other approaches [11].